OpenStudy (anonymous):

Find the sum 100 ∑ (3 + k) k=1

OpenStudy (anonymous):

\[\sum_{k=1}^{100}3=3\times 100\] \[\sum_{k=1}^{100}k=\frac{100\times 101}{2}=5050\]

OpenStudy (anonymous):

add to get \[5350\]